Find an equation of the line satisfying the given condtions.
Through (0,2); m=3/2
possible answers:
2x-3y=-4
3x+2y=-4
3x-2y=4
3x-2y=-4

Answer by Alan3354(69443)

Through (0,2); m=3/2
------------
y-y1 = m*(x-x1) where (x1,y1) is the point
y-2 = (3/2)x
2y-4 = 3x
3x-2y = -4